What is the minimum work needed to push a 800 -kg car 820 m up along a 9.5 ∘ incline? Ignore friction.

Express your answer to two significant figures and include the appropriate units.
Answer #1

* 820 0-9.50 Sin 9.5= h 820 h=820x81 19.57 135-33m W = mgh = 800×9.8x 135-331061058.04] Cs Scanned with CamScanner